OpenStudy (anonymous):

A U.S. football field is 139.5 feet longer than it is wide. If its perimeter is 921 feet, find its demensions.

OpenStudy (anonymous):

921 = (139.5 ft + W) + W 921 = 139.5 ft + 2W 921 - 139.5 ft = 2W 781.5 ft = 2W 390.75 = W to check: 921 = (139.5 + 390.75) + 390.75 921 = 530.25 + 390.75 921 = 921 or 530.25 - 390.75 = 139.5 longer than the width :) is it right ????.???

OpenStudy (anonymous):

I believe so thank you soo much for your help

OpenStudy (anonymous):

your welcome :)